# Learning on the Fly: Rapid Policy Adaptation via Differentiable Simulation

### v1

@@ IV-A2 Pretraining Phase
We parameterize the policy as an MLP with two 512-dim hidden layers. For both state-based hovering and trajectory tracking, we train the base policy from random initialization for 300 epochs across 100 parallel environments. Each epoch lasts 3 seconds or 150 simulation steps. For visual feature-based hovering, we use the initialization approach from [20] to first train a neural network on a state-representation learning task and use the learned parameters to partially initialize the policy network. This has been shown to improve convergence and sample efficiency in policy learning [39]. We refer the reader to [20] for details on this initialization method. We then train the partially initialized policy for 500 epochs across 300 parallel environments.
@@ IV-A3 Online Adaptation Phase
The quadrotor states and actions are continuously recorded into a rolling history buffer at 50 and are used to train the residual dynamics network. For stabilizing hover and trajectory tracking, we use history buffer sizes of 100 and 250, equivalent to 2 and 5 seconds of past quadrotor states and actions, respectively. For residual dynamics learning, we continuously refine an ensemble of 3 networks, each with two 128-dim hidden layers and initialized using different random seeds, and use the empirical mean prediction from all models as the final predicted residual acceleration for a given input. Empirically, we found this to effectively reduce the prediction variance arising from epistemic uncertainty due to the limited samples in the data buffer. We run residual dynamics learning every 3 seconds and train the ensemble networks in parallel for 100 iterations. Policy adaptation is run every 5 seconds, and we train the state-based policy with 10 parallel simulated environments for 30 epochs and the vision-based policy with 30 environments for 50 epochs. These values were empirically found to provide a good balance between training time and policy performance.
@@ IV-A4 Baselines Methods
We compare against a state-of-the-art learning-based adaptive control method, Deep Adaptive Tracking Control (DATT) [13], which uses the popular model-free RL algorithm PPO with domain randomization and online adaptive control-based disturbance estimation. For quadrotor control, this method has been shown to outperform Rapid Motor Adaptation (RMA) [4], which is a similar approach but instead uses a learned encoder for disturbance estimation. We used the open-source implementation of [13] and the exact same training procedure and hyperparameters to train both state-based hovering and trajectory tracking policies using PPO for 20 million simulation steps. Using their original domain randomization method, we simulated 3-dimensional acceleration disturbances as random walks within the bounds +/- . We also compare against an adaptive Nonlinear MPC controller (-MPC) as implemented in [13], which uses a Model Predictive Path Integral (MPPI) formulation and the same adaptive control-based disturbance estimation as in DATT. Additionally, we include our pretrained base policy (Base DiffSim) for each task without further online adaptation for comparison.
@@ IV-B Experimental Results
For evaluation, we used a realistic quadrotor simulator [12] equipped with the BEM model for aerodynamic effects and high-frequency simulations of the controller dynamics. We simulated three levels of constant, uniform acceleration disturbances: (none), (small), and (large). The first two conditions are within the domain randomization range used for DATT training, whereas the third condition was deliberately chosen to be out-of-distribution to evaluate its generalization capabilities. All experiments (both simulated and real-world) were run using an Nvidia RTX 4090 GPU with an Intel 14900KF CPU.
@@ IV-B1 Performance Comparison to Baseline Approaches
For the state-based stabilizing hover task, we ran each method under all disturbance conditions from a set of 8 different starting positions around the hovering target, and used the final steady-state error as the performance metric. To ensure a fair comparison, we continued running each method until no further accuracy improvements were observed. This was found to be approximately 10 seconds for all baseline methods as they do not require any policy adaptation, and around 30 seconds for our approach (both state and vision-based) for a few learning steps to take place. As summarized in Tab. I, results show that our method consistently exhibits superior or comparable performance to the baselines. Fig. 5(a) illustrates that our method rapidly adapts the policy to compensate for the large disturbances within 2-3 adaptation steps. DATT performs well under both the none and small disturbance scenarios which are within its training distribution, but struggles to adapt to the larger, out-of-distribution disturbance.
@@
| Base DiffSim | 0.365 0.124 | 0.571 0.091 | 1.479 0.213 | |
| Circle | -MPC | 0.113 0.027 | 0.096 0.063 | 0.410 0.155 |
| DATT (PPO) | 0.058 0.016 | 0.040 0.024 | crash | |
| Ours | 0.167 0.048 | 0.135 0.101 | 0.349 0.175 | |
| Ours (LoRA) | 0.129 0.051 | 0.159 0.043 | 0.326 0.061 | |
@@
| Ours | 0.068 0.040 | 0.045 0.03 | 0.137 0.098 | |
| Ours (LoRA) | 0.069 0.047 | 0.059 0.043 | 0.110 0.037 |
For trajectory tracking, we recorded 60-second rollouts and computed the average tracking error () within the last 10-second window as the performance metric. As shown in Tab. II, our approach achieves comparable performance to the baselines for both Circle and Figure-8 across all disturbance conditions. We observed that our method is able to rapidly adapt and achieve much improved tracking accuracy after only 3-4 policy update steps. For DATT, consistent with findings from state-based hovering, it fails to generalize to the out-of-distribution disturbances and results in crashes for both reference trajectories.
@@
| Ours (State) | 0.015 0.001 | 0.008 0.002 | 0.105 0.007 |
| Ours (Vision) | 0.082 0.009 | 0.099 0.021 | 0.207 0.041 |
| Ours (Vision, LoRA) | 0.084 0.014 | 0.111 0.024 | 0.205 0.048 |
@@ IV-B2 Computational and Sample Efficiency Analysis
We analyze and compare the sample and computational efficiency of our approach against DATT for state-based tasks. For our approach, policy pretraining uses 300 epochs across 100 environments, which is equivalent to 4.5 million simulation steps in total, and takes approximately 15 seconds. Empirically, we observed that good-performing initial policies can in fact be obtained using fewer epochs and environments, thanks to the low-variance first-order policy gradients from differentiable simulation. For online adaptation, each residual dynamics learning step (100 iterations) takes approximately 2 seconds, and each policy adaptation step runs for 30 epochs across 10 environments (or 45k simulation steps) and takes about 1.5 seconds. Here, with only 3 adaptation steps, which is equivalent to 4.5 seconds of policy training in wall time, we already observe significant performance improvements for both hovering and tracking. In comparison, DATT trains the policy for 20 million simulation steps, which takes around 2 hours, and requires no further training at runtime. For DATT, we also observed slower convergence to lower rewards when training with larger domain randomization, which is likely a result of the performance-generalization trade-off [40], possibly exacerbating the high variance in the policy gradient estimates. In summary, our approach enables more efficient compute usage by simplifying initial policy training without domain randomization or curricula, and by supporting sample- and compute-efficient online adaptation to out-of-distribution scenarios where domain randomization fails to generalize.
@@ IV-B3 Continuous Adaptation to Time-Varying Disturbances
We demonstrate the ability of our method to continuously adapt policies to unknown time-varying disturbances, using the realistic quadrotor simulator [12]. Here, we show an example of continuously adapting a state-based hovering policy under uniform, time-varying acceleration disturbance given by , which reverses its direction every 15 . As shown in Fig. 5(b), our approach rapidly adapts the state-based policy within 2 adaptation steps to adjust for each disturbance change, with the policy behavior remaining stable throughout the entire process. We observed the continuous adaptation of a visual feature-based hovering policy to be less stable than its state-based counterpart, consistent with our previous findings. For continuous trajectory tracking adaptation experiments, we provide detailed visualizations in the supplementary material.
@@ IV-C Optimizing Runtime Efficiency and Performance
We conducted an analysis using the state-based stabilizing hover task to justify two key design choices in the differentiable simulation pipeline: 1) low-fidelity analytical dynamics model for simulation, and 2) gradient backpropagation only through the analytical dynamics model. Given that a key objective is to minimize runtime while maintaining policy performance, we compared the training time and policy performance for each design choice. We used the same realistic quadrotor simulator [12] and added a constant uniform acceleration disturbance of 2 in the positive x-axis direction. We first collected 50 rollout trajectories (3 seconds each) of the base hovering policy from 50 random starting configurations around the hovering target, and then used the generated residual samples to train a single residual dynamics network for 200 epochs. Finally, we adapted the base policy by running 100 epochs of policy training across 100 parallel environments, and evaluated the final steady-state errors from the hovering target across 8 rollout trajectories.
We first compared using the low-fidelity (low-fid) dynamics model (1) against using a high-fidelity model (high-fid) which simulates body and rotor drag effects, the rotor thrust maps, and the low-level controller dynamics, as the analytical dynamics model together with the residual dynamics network for forward simulation (res-fwd). We found that using the low-fidelity model achieves approximately 2-times faster training (see Fig. 6(a)) than using the high-fidelity model, while the achieved policy performances by both methods were very comparable (see Fig. 6(b)). For gradient backpropagation, we found that in addition to backpropagating through the analytical dynamics model, also performing backpropagation through the learned residual dynamics network (res-back) increases training time by approximately 30% without providing clear benefits to the policy performance. This is consistent with previous findings [20, 11] that combining accurate forward simulation with a surrogate gradient which points in approximately the same direction as the true gradient vector accelerates policy training without impacting the resulting policy performance.
@@ IV-D Real World Validation
We validated our approach in real-world experiments for both stabilizing hover and trajectory tracking tasks using the exact same procedures as the simulated experiments for both pretraining and online adaptation. Detailed visualizations and experiments are provided in the supplementary material. A motion-capture system provides state estimation of the quadrotor at 100 to the off-board workstation, which computes and send commands to the on-board flight controller at 50 . To create some variation in the real-world dynamics, we used two quadrotors adapted from the Agilicious platform [12]: a small lightweight quadrotor and a larger, heavier one with different dynamical properties (see Tab. IV). Moreover, we modified the small quadrotor by rigidly attaching to it a quadrotor stand from below, increasing its mass from 190 to 260 by approximately while simultaneously altering its inertial properties. Finally, we used a fan to create wind disturbances, resulting in complex, state-dependent forces on the modified quadrotor due to its highly imbalanced and non-uniform drag profile. Therefore, both the existing sim-to-real gap and the extra disturbances contribute to significant out-of-distribution dynamics that were unseen during policy pretraining.
Fig. 7(a) shows the recorded trajectory of adaptation of the state-based hovering policy on the modified small quadrotor under a diagonal wind disturbance. Despite the more complex and unstable real-world disturbance forces compared to the constant uniform disturbance in simulation, our method still enables the policy to rapidly adapt with 2-3 policy update steps to compensate for disturbances. Similar results were observed for visual feature-based hovering, where the adaptation process appeared less stable than state-based hovering, which is consistent with our findings from the simulated experiments. Real-world experiments also show that our approach achieves accurate trajectory tracking under significant model mismatches. Here, we present one particular experiment where a base Figure-8-tracking policy was deployed on the large quadrotor. As shown in Fig. 7(b), despite the poor initial tracking and state-space exploration caused by the large sim-to-real gap, the policy quickly adapts and achieves much improved tracking within just a few policy update steps.
@@ V Conclusion
We propose a novel rapid policy adaptation framework combining online residual dynamics learning from real-world flight data and sample-efficient policy learning via differentiable simulation. With all system components designed for rapid adaptation, we demonstrate the possibility to adapt both state and visual feature-based policies to unknown disturbances within several seconds. One limitation of our framework lies in the tightly-coupled dependencies between data collection via policy rollout and policy learning using learned residual dynamics from the collected data. The quality and rate of convergence may be affected by biases or noise in the learned residual dynamics. The dependency is closely related to the concept of performative prediction in related machine learning fields. Thus, future work will explore uncertainty-driven data collection where the policy is augmented by active exploration to simultaneously improve task performance and reduce uncertainty in the real-world dynamics.

### v2

@@ IV-A2 Pretraining Phase
We parameterize the policy as an MLP with two 512-dim hidden layers. For both state-based hovering and trajectory tracking, we train the base policy from random initialization for 300 epochs across 100 parallel environments. Each epoch lasts 3 seconds or 150 simulation steps. For visual feature-based hovering, we use the initialization approach from [heeg2024learning] to first train a neural network on a state-representation learning task and use the learned parameters to partially initialize the policy network. We refer the reader to [heeg2024learning] for details on this initialization method. We then train the partially initialized policy for 500 epochs across 300 parallel environments.
@@ IV-A3 Online Adaptation Phase
The quadrotor states and actions are continuously recorded into a rolling history buffer at 50 and are used to train the residual dynamics network. For stabilizing hover and trajectory tracking, we use history buffer sizes of 100 and 250, equivalent to 2 and 5 seconds of trajectory history, respectively. For residual dynamics learning, we continuously refine an ensemble of 3 networks, each with two 128-dim hidden layers and initialized using different random seeds, and use the empirical mean prediction from all models as the final predicted residual acceleration for a given input. Empirically, we found this to effectively reduce the prediction variance arising from epistemic uncertainty due to the limited samples in the data buffer. For LoRA, we follow the original implementation and initialization in [hu2022lora], and use ranks of 4, 10, and 1 for the three weight matrix layers of the policy network, respectively, and a constant scale of 4 across all layers. We run residual dynamics learning every 3 seconds and train the ensemble networks in parallel for 100 iterations. Policy adaptation is run every 5 seconds, and we train the state-based policy with 10 parallel simulated environments for 30 epochs and the vision-based policy with 30 environments for 50 epochs. These values were empirically found to provide a good balance between training time and policy performance.
@@ IV-A4 Baselines Methods
We compare against a state-of-the-art learning-based adaptive control method, Deep Adaptive Tracking Control (DATT) [huang2023datt], which uses the popular model-free RL algorithm PPO with domain randomization and online adaptive control-based disturbance estimation. For quadrotor control, this method has been shown to outperform Rapid Motor Adaptation (RMA) [kumar2021rma], which is a similar approach but instead uses a learned encoder for disturbance estimation. We used the open-source implementation of [huang2023datt] and the exact same training procedure and hyperparameters to train both state-based hovering and trajectory tracking policies using PPO for 20 million simulation steps. Using their original domain randomization method, we simulated 3-dimensional acceleration disturbances as random walks within the bounds +/- . We also compare against an adaptive Nonlinear MPC controller (-MPC) as implemented in [huang2023datt], which uses a Model Predictive Path Integral (MPPI) formulation and the same adaptive control-based disturbance estimation as in DATT. Additionally, we include our pretrained base policy (Base DiffSim) without online adaptation for comparison.
@@ IV-B Experimental Results
We used a realistic quadrotor simulator [foehn2022agilicious] equipped with the BEM model for aerodynamic effects and high-frequency simulation of controller dynamics. We simulated three levels of constant, uniform acceleration disturbances: (none), (small), and (large). The first two conditions are within the domain randomization range used for DATT training, whereas the third condition was deliberately chosen to be out of distribution to evaluate its generalization capabilities. All experiments (simulated and real-world) were run using an Nvidia RTX 4090 GPU (24 GB VRAM) with an Intel 14900KF CPU.
@@ IV-B1 Performance Comparison to Baseline Approaches
For the state-based stabilizing hover task, we ran each method under all disturbance conditions from a set of 8 different starting positions around the hovering target, and used the final steady-state error as the performance metric. To ensure a fair comparison, we continued running each method until no further accuracy improvements were observed. This was found to be approximately 10 seconds for all baseline methods, as they do not require any policy adaptation, and around 30 seconds for our approach (both state and vision-based) for a few learning steps to take place. As summarized in Tab. I, results show that our method consistently exhibits superior or comparable performance to the baselines. Fig. 5(a) illustrates that our method rapidly adapts the policy to compensate for the large disturbances within 2-3 adaptation steps. DATT performs well under both the none and small disturbance scenarios, which are within its training distribution, but struggles to adapt to the larger, out-of-distribution disturbance.
@@
| Circle | Base DiffSim | 0.365 0.124 | 0.571 0.091 | 1.479 0.213 |
[…]
@@
[…]
| 5-Point Star | Base DiffSim | 0.453 0.190 | 0.467 0.236 | 0.844 0.389 |
| -MPC | 0.295 0.086 | 0.218 0.111 | 0.417 0.086 | |
[…]
For trajectory tracking, we recorded 60-second rollouts and computed the average tracking error () within the last 10-second window as the performance metric. As shown in Tab. II, our approach achieves comparable performance to the baselines across all trajectories and disturbance conditions. Here, our method exhibits consistent responsiveness to both smooth and non-smooth references. In particular, for the 5-Point Star, our method achieves comparable tracking accuracy as DATT, which has demonstrated strong capabilities in tracking non-smooth, dynamically infeasible trajectories [huang2023datt]. We observed that our method is able to rapidly adapt and achieve much improved tracking accuracy after only 3-4 policy update steps. For DATT, consistent with findings from state-based hovering, it fails to generalize to the out-of-distribution disturbances and results in crashes for all reference trajectories.
@@
[…]
@@ IV-B2 Computational and Sample Efficiency Analysis
We analyze and compare the sample and computational efficiency of our approach against DATT for state-based tasks. For our approach, policy pretraining uses 300 epochs across 100 environments, which is equivalent to 4.5 million simulation steps in total, and takes approximately 15 seconds. Empirically, we observed that good-performing initial policies can in fact be obtained using fewer epochs and environments, thanks to the low-variance first-order policy gradients from differentiable simulation. For online adaptation, each residual dynamics learning step (100 iterations) takes approximately 2 seconds, and each policy adaptation step runs for 30 epochs across 10 environments (or 45k simulation steps) and takes about 1.5 seconds. Here, with only 3 adaptation steps, which is equivalent to 4.5 seconds of policy training in wall time, we already observe significant performance improvements for both hovering and tracking. In comparison, DATT trains the policy for 20 million simulation steps, which takes around 2 hours, and requires no further training at runtime. For DATT, we also observed slower convergence to lower rewards when training with larger domain randomization, which is likely a result of the performance-generalization trade-off [zhou2022domain], possibly exacerbating the high variance in the policy gradient estimates. In summary, our approach enables more efficient compute usage by simplifying initial policy training without domain randomization or curricula, and by supporting sample- and compute-efficient online adaptation to out-of-distribution scenarios where domain randomization fails to generalize.
@@ IV-B3 Continuous Adaptation to Time-Varying Disturbances
We demonstrate the ability of our method to continuously adapt policies to unknown time-varying disturbances, using the realistic quadrotor simulator [foehn2022agilicious]. Here, we show an example of continuously adapting a state-based hovering policy under uniform, time-varying acceleration disturbance given by , which reverses its direction every 15 . As shown in Fig. 5(b), our approach rapidly adapts the state-based policy within 2 adaptation steps to adjust for each disturbance change, with the policy behavior remaining stable throughout the entire process. We observed that the continuous adaptation of a visual feature-based hovering policy was less stable than its state-based counterpart, consistent with our previous findings. For continuous trajectory tracking adaptation experiments, we provide detailed visualizations in the supplementary material.
@@ IV-C Optimizing Runtime Efficiency and Performance
We conducted an analysis using the state-based stabilizing hover task to justify two key design choices in the differentiable simulation pipeline: 1) low-fidelity analytical dynamics model for simulation, and 2) gradient backpropagation only through the analytical dynamics model. Given that a key objective is to minimize runtime while maintaining policy performance, we compared the training time and policy performance for each design choice. We used the same realistic quadrotor simulator [foehn2022agilicious] and added a constant uniform acceleration disturbance of 2 in the positive x-axis direction. We first collected 50 3-second rollout trajectories of the base hovering policy from random starting positions around the target, and then used the generated residual samples to train a single residual dynamics network for 200 epochs. Finally, we adapted the base policy by running 100 epochs of policy training across 100 parallel environments, and evaluated the final steady-state errors from the hovering target across 8 rollout trajectories.
We first compared using the low-fidelity (low-fid) dynamics model (1) against using a high-fidelity model (high-fid), which simulates body and rotor drag effects, rotor thrust maps, and low-level controller dynamics, as the analytical dynamics model together with the residual dynamics network for forward simulation (res-fwd). We found that using the low-fidelity model achieves approximately 2-times faster training (see Fig. 6(a)) than using the high-fidelity model, while the achieved policy performances by both methods were very comparable (see Fig. 6(b)). For gradient backpropagation, we found that in addition to backpropagating through the analytical dynamics model, also performing backpropagation through the learned residual dynamics network (res-back) increases training time by approximately 30% without providing clear benefits to the policy performance. This is consistent with previous findings [heeg2024learning, song2024learning] that combining accurate forward simulation with a surrogate gradient that points in approximately the same direction as the true gradient vector accelerates policy training without impacting the resulting policy performance.
@@ IV-D Real World Validation
We conducted real-world experiments using the same tasks as the simulated experiments and the same pretraining and online adaptation procedures. A motion-capture system provides quadrotor state estimation at 100 to the off-board workstation, which computes and sends commands to the on-board controller at 50 . We used two quadrotors adapted from the Agilicious platform [foehn2022agilicious]: a small, lightweight quadrotor and a larger, heavier one with different dynamical properties (see Tab. IV). Moreover, we modified the small quadrotor by rigidly attaching to it a quadrotor stand from below, increasing its mass from 190 to 260 by approximately and altering its inertial properties. Finally, we used a fan to create wind disturbances, resulting in complex, state-dependent forces on the modified quadrotor due to its highly imbalanced and non-uniform drag profile. Both the existing sim-to-real gap and the extra disturbances contribute to significant out-of-distribution dynamics that were unseen during policy pretraining.
Fig. 7(a) shows state-based hovering adaptation on the modified small quadrotor under a diagonal wind disturbance. Despite the more complex and unstable real-world disturbance forces compared to the constant uniform disturbance in simulation, our method still enables the policy to rapidly adapt with 2-3 policy update steps to compensate for disturbances. Similar results were observed for visual feature-based hovering, where the adaptation process appeared less stable than state-based hovering, which is consistent with our findings from simulated experiments. Real-world experiments also show that our approach achieves accurate trajectory tracking under added mass, wind, and significant model mismatches. In particular, our method achieves accurate tracking of the non-smooth 5-Point Star under complex wind disturbances (see Fig. 7(b)). Moreover, Fig. 7(c) shows one particular experiment where a Figure-8-tracking policy was deployed on the large quadrotor. Despite the poor initial tracking and state-space exploration caused by the large sim-to-real gap, the policy quickly adapts and achieves much improved tracking within just a few policy update steps. Videos and visualizations of real-world experiments are provided in the supplementary material.
@@ V Conclusion
We propose a novel rapid policy adaptation framework combining online residual dynamics learning from real-world flight data and sample-efficient policy learning via differentiable simulation. With all system components designed for rapid adaptation, we demonstrate the possibility to adapt both state and visual feature-based policies to unknown disturbances within several seconds. One limitation of our framework lies in the tightly-coupled dependencies between data collection via policy rollout and policy learning using learned residual dynamics from the collected data. The quality and rate of convergence may be affected by biases or noise in the learned residual dynamics. Thus, future work will explore uncertainty-driven data collection where the policy is augmented by active exploration to simultaneously improve task performance and reduce uncertainty in the real-world dynamics.